OpenStudy (anonymous):

Write a linear system that models the situation. The sum of two numbers is 14. Their difference is 10. I got 2+12 = 14. Help? I suck at this...

OpenStudy (anonymous):

Okay I'm stumped.

OpenStudy (anonymous):

Does the problem call for you to solve it?

OpenStudy (anonymous):

I don't think so. It just says model the situation as a linear equation..

OpenStudy (anonymous):

Maybe it's just x + y = 14

OpenStudy (anonymous):

As a linear "system"?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

A system of equations has more than one equation.

OpenStudy (anonymous):

That's why I don't understand this...It's weird.

OpenStudy (anonymous):

The term linear refers to the variables having no exponents. So, it's asking you to write those two equations using x and y

OpenStudy (anonymous):

Don't you mean no coefficients?

OpenStudy (anonymous):

Your first answer, X + Y = 14 is correct. You need to turn the Something minus something = 10 into a equation

OpenStudy (anonymous):

Oh.

OpenStudy (anonymous):

Uh x -y = 10

OpenStudy (anonymous):

Bam. That's it! X + Y = 14 X - Y = 10

OpenStudy (anonymous):

Wow...

OpenStudy (anonymous):

And yup, a linear equation has no exponents. All of it's variables are to the first power.

OpenStudy (anonymous):

Damn I'm stupid

OpenStudy (anonymous):

I always think the problem is going to be hard, but it's hella easy.

OpenStudy (anonymous):

Thank You once...again.

OpenStudy (anonymous):

You're welcome. Also, Don't ever call yourself stupid, because you're not. Don't sell yourself short. If you keep telling yourself your stupid, you'll eventually believe it.

OpenStudy (anonymous):

I'm stupid on this one question.

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

Okay now I'm stuck again.

OpenStudy (anonymous):

What is an equation for one CD costing $2 more than the other? Do you know? >-<

OpenStudy (anonymous):

Well, the first step is to assign variables, right?

OpenStudy (anonymous):

You told me CDA = CD2 + $2, right?

OpenStudy (anonymous):

Well both of the two Cds equal $28

OpenStudy (anonymous):

We can use mathematics to describe relationships between things

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

What's the whole question?

OpenStudy (anonymous):

You buy two CDs for a total of $28. One CD cost $2 more than the other.

OpenStudy (anonymous):

-_- XC Bahh. I can't answer it.

OpenStudy (anonymous):

? Sure you can.

OpenStudy (anonymous):

It's asking you to write another system of equations, right?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

First off, what's in a system of equations? What is it? And why do we use it?

OpenStudy (anonymous):

We try to find the values of the unknown?

OpenStudy (anonymous):

A system of equations is a list of equations. It has several equations, usually with more than one variable

OpenStudy (anonymous):

Are you still here?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Okay. We use a system of equations to help solve for more than one variable.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

With one equation, X + Y = 12, it's impossible to find what X and Y are. Can you tell me why?

OpenStudy (anonymous):

Please tell me why. ^^

OpenStudy (anonymous):

Because we need a real number. You can't just solve it with variables?

OpenStudy (anonymous):

Nope. Try graphing it

OpenStudy (anonymous):

You can't.

OpenStudy (anonymous):

unless you make up the numbers..

OpenStudy (anonymous):

Variables exist within an equation, so we can put numbers in.

OpenStudy (anonymous):

I think what you mean, is that there is an infinite amount of solutions, right?

OpenStudy (anonymous):

Yeah..pretty much.

OpenStudy (anonymous):

In the equation (X+Y = 12), if X = 3, than Y = 9, if X is 10, than Y is 2, and so on

OpenStudy (anonymous):

You can graph these points, and it makes a straight line. The straight line represents all of the answers that work in the equation.

OpenStudy (anonymous):

Okay. Are you telling me then I can find the slope...? rise/run right?

OpenStudy (anonymous):

The slope is a tool. It helps us graph the line, and it tells us how quickly it goes up or down. I'm explaining this to you, because I don't think you completely grasp graphing, or equations yet.

OpenStudy (anonymous):

I know how to graph equations.

OpenStudy (anonymous):

I usually use y=mx + b

OpenStudy (anonymous):

I then graph it from there, or there are other forms I can graph from like standard form or point-slope form.

OpenStudy (anonymous):

I undestand that you can graph. But do you understand why you graph?

OpenStudy (anonymous):

Why do we graph?

OpenStudy (anonymous):

We graph for a couple of reasons.

OpenStudy (anonymous):

First off, pictures are easier to understand than really complicated formulas

OpenStudy (anonymous):

True,

OpenStudy (anonymous):

Next, with a graph, we can solve equations that we don't know the answer of

OpenStudy (anonymous):

Equations? That's only useful for math! But, remember what I said? We use math to describe how things are related

OpenStudy (anonymous):

I'm going to give you two equations, and I want you to graph them, okay? When you graph them, you should be able to get the answer without doing any algebra.

OpenStudy (anonymous):

You need to just plug in a few points, and then plot them.

OpenStudy (anonymous):

X + Y = 12, and X - Y = 5

OpenStudy (anonymous):

Graph them on the same graph.

OpenStudy (anonymous):

(If that makes any sense :) )

OpenStudy (anonymous):

Yeah it does. You can solve linear systems by substitution or graphing or elimination..

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

Okay it's 2:08am. It's not hat I want to not learn, but I need to goto sleep now. Thanks for the lecture.

OpenStudy (anonymous):

If you graph the two lines, they should intersect at only one point. That's the only answer that's correct.

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

For your question, you need to write a system of equations.

OpenStudy (anonymous):

Yeah they intersect, but I still think substitution is the easiest way out of all the methods I've learned.

OpenStudy (anonymous):

It's the fastest. With the more complicated math, you'll suddenly love graphing

OpenStudy (anonymous):

What itf you can't draw lines that well?

OpenStudy (anonymous):

Erm, substitution is. And, the lines don't have to be perect.Anyway, You buy two CDs for a total of $28. One CD cost $2 more than the other.

OpenStudy (anonymous):

also I know I have to write linear systems but I only think I have to write two.

OpenStudy (anonymous):

x + y = 28 I know that's one.

OpenStudy (anonymous):

I think the other is 2 - y ...

OpenStudy (anonymous):

One CD + 2 = another CD, right?

OpenStudy (anonymous):

Just substitute the variables

OpenStudy (anonymous):

No the total cost of the two equals 28 dollars.

OpenStudy (anonymous):

I don't need to solve it..

OpenStudy (anonymous):

Yeah. But the 2nd clause says that one cd costs two more dollars than the other

OpenStudy (anonymous):

That's an equation.

OpenStudy (anonymous):

Oh I wrote an expression..

OpenStudy (anonymous):

Oh I wrote an expression..

OpenStudy (anonymous):

Oh I wrote an expression..

OpenStudy (anonymous):

?

OpenStudy (anonymous):

X + 2 = Y. Truthfully, it doesn't matter what variable +2 = the other variable.

OpenStudy (anonymous):

Must be a glitch

OpenStudy (anonymous):

Anyway, I'm sorry I wasn't the most helpful for the equation. God bless with your mathematic endeavors, and good night.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Ashag, are you still here? If so, go to your latest problem, and I"ll take a crack at it